# Rank-1 Approximation of Inverse Fisher for Natural Policy Gradients in Deep Reinforcement Learning

### v1

@@ 4.3 SM-ActorCritic
using the Adam optimizer, where $R_{t}=A_{t}+V_{\phi}(s_{t})$ is the Monte-Carlo return target. The actor is updated by a single natural policy gradient step obtained from the matrix-free Fisher inverse. The full SMAC update is given in Algorithm 1.
@@
7:   $g^{k}\leftarrow\nabla_{\theta}\!\log\pi_{\theta}(a\!\mid\!s)$
@@
10:  update critic: $\phi^{k+1}\leftarrow\phi^{k}-\alpha\nabla_{\phi}\sum_{t}(R^{k}-V_{\phi})^{2}$
@@ 5 Experimental Setup
To evaluate the performance of our proposed approach, SMAC, we provide experiments in several environments and compare against three baselines. We begin with three simple control environments, before moving on to six MuJoCo tasks (Todorov et al., 2012). For all experiments, we used the Gymnasium framework (Towers et al., 2024).
We implement SMAC on top of the A2C algorithm. In all environments, we evaluate SMAC against three baselines. To measure the benefits of using our NPG approximator, we consider an algorithm that is similar to SMAC in all aspects, except that it uses first-order updates instead of natural gradients. This is equivalent to A2C with a simple stochastic gradient descent (SGD) optimizer, which we denote as AC-SGD. To ensure a fair comparison to vanilla A2C, we also compare against the AC-Adam baseline, where A2C is trained using the more complex Adam optimizer. Finally, we also test SMAC against a different NPG approximator. The AC-CG baseline uses the conjugate gradient algorithm to directly compute the NPG during the A2C update, without ever computing or storing the Fisher information matrix itself, similarly to TRPO (Schulman et al., 2015a). To better handle continuous control problems, SMAC and all the baselines we compare against compute advantages using the GAE (Schulman et al., 2015b).
For each of the four algorithms in each environment, we train five models using different random seeds. The results reported throughout training are averaged over these seeds, with standard deviations included. We evaluate agent performance using (undiscounted) returns, averaged across episodes. Besides the maximum average returns achieved, we are also interested in convergence speed. Furthermore, we discuss the average log-probabilities of the actions taken as a measure of the agents’ confidence in the actions they take. We report these in detail in the supplementary material.
In all environments, episodes have a horizon of $1000$ timesteps. To increase the readability of the plots, we group timesteps into $100$ bins with no overlap and report the average of each bin. We apply additional smoothing by computing an exponentially weighted mean over the bins, with a smoothing factor of $0.1$.
AC-SGD AC-Adam AC-CG SMAC (Ours)
Acrobot $\mathbf{-86.2\pm 5.0}$ $88.4\pm 3.8$ $93.3\pm 20.3$ $94.1\pm 6.3$
Cartpole $962.3\pm 49.0$ $\mathbf{989.5\pm 0.3}$ $977.9\pm 0.0$ $973.4\pm 29.8$
Pendulum $2116\pm 564$ $5731\pm 150$ $\mathbf{-190\pm 18}$ $2226\pm 859$
Half Cheetah $888.7\pm 15.4$ $746.3\pm 24.6$ $444.6\pm 12.1$ $\mathbf{863.8\pm 38.7}$
Hopper $180.5\pm 13.6$ $303.8\pm 27.1$ $172.7\pm 16.3$ $\mathbf{331.2\pm 29.7}$
Swimmer $12.4\pm 5.2$ $17.3\pm 0.5$ $23.8\pm 0.9$ $\mathbf{28.5\pm 0.8}$
Walker $213.4\pm 42.9$ $79.3\pm 95.0$ $\mathbf{229.8\pm 23.9}$ $204.1\pm 87.9$
Humanoid $4539\pm 110$ $3105\pm 299$ $3175\pm 1423$ $\mathbf{4625\pm 277}$
Pusher $433.6\pm 54.9$ $568.3\pm 12.8$ $441.0\pm 34.5$ $\mathbf{-408.2\pm 32.2}$
Table 1: Average returns per episode, measured at the end of training, for each task in both classic control and MuJoCo environments. All results are averaged over 55 seeds.
5.1 Ablation: Effect of Batch Size
The matrix-free updates operate on a single trajectory sample. Although the per-step cost is only $\mathcal{O}(d)$, many more updates are required to process a fixed number of transitions, so the computational cost can be very high.
To investigate this trade-off, we evaluate an alternative that uses the average gradient of log-probability:
@@ 5.1 Ablation: Effect of Batch Size
The regularized empirical Fisher then becomes
${\hat{F}^{k}}=\lambda\bm{I}+\bar{\ell}^{k}\,(\bar{\ell}^{k})^{\top},$
@@
$\Delta\theta^{k}=\eta\left[\frac{1}{\lambda}g^{k}-\frac{\bar{\ell}^{k}\,(\bar{\ell}^{k})^{\top}g^{k}}{\lambda^{2}+\lambda(\bar{\ell}^{k})^{\top}\bar{\ell}^{k}}\right].$
@@ 6 Results
We provide the final performance achieved by training each of the four algorithms on each of the nine tasks. These results, presented in Table 1, show that our method, SMAC, can achieve competitive results in two of the three simple control tasks. More importantly, the benefits of SMAC appear to increase in more complex environments, such as MuJoCo, where SMAC outperforms the other baselines in terms of final returns in all but one task. A more in-depth analysis of these results, with a special focus on convergence rate, is given in the following sections.
Figure 2: Results on three classic control tasks. We report the average return per episode. All results are averaged over 55 random seeds, with shaded areas showing standard deviation.
@@ 6.1 Classic Control
In the classic control environment, our method, SMAC, reaches a high initial average return faster than the baselines on two of the tasks, as shown in Fig. 2. This, however, comes at the cost of additional variance across seeds. The overall performance of SMAC is also competitive in those two environments.
In Acrobot, SMAC only requires $40000$ timesteps to pass the average return threshold of 400-400. This is faster than AC-Adam, AC-SGD, and AC-CG, which require $70\%$, $35\%$, and $50\%$ more timesteps, respectively. This accelerated learning is, however, followed by unstable training until approximately $125000$ timesteps. Despite this, SMAC stabilizes and converges, reaching a final performance of $94.1\pm 6.3$, which is similar to the baselines.
A similar behavior can be observed in Cartpole. SMAC agents achieve $75\%$ of the maximum return possible in $87000$ timesteps, slightly faster than AC-CG, which needs approximately $10\%$ more timesteps. AC-Adam and AC-SGD are even slower, requiring $28\%$ and $62\%$ more timesteps, respectively, to reach the same threshold. SMAC becomes more unstable after this initial period, but stabilizes, reaching a competitive performance of $973.4\pm 29.8$ by the end of training.
In contrast, in the last environment, Pendulum, our method fails to outperform AC-CG in terms of convergence speed and overall performance. AC-CG requires only $1.5$ million timesteps to match the performance of SMAC, indicating that our proposed approximation of the FIM can struggle to improve learning in some environments. However, SMAC still achieves a final return of $2226\pm 859$, outperforming AC-Adam and matching AC-SGD, with the former also displaying much slower and unstable learning.
The validity of our results is also confirmed by the action log-probabilities recorded. In Acrobot and Cartpole, the two tasks in which SMAC performs competitively, the log-probabilities converge faster than those of the baseline methods. Together with the increase in performance achieved by SMAC in these tasks, this may indicate that our agents can rapidly become confident in their policies when taking actions that yield increasingly higher returns. At the same time, in Pendulum, the log-probabilities of SMAC match those of AC-SGD and are bounded above by AC-CG. This could indicate that our method becomes overconfident in its actions too early, finally leading to a sub-optimal policy.
@@ 6.2 MuJoCo
In the more complex MuJoCo environment, we have found that throughout training, our method, SMAC, consistently outperforms or matches the baseline algorithms in terms of performance in $4$ out of $6$ tasks (Fig. 3). In the other two, SMAC still achieves comparable results. Furthermore, SMAC exhibits a noticeably higher sample efficiency in three tasks.
@@
Figure 3: Results on six MuJoCo tasks. We report the average return per episode. All results are averaged over 55 random seeds, with shaded areas showing standard deviation.
The largest improvement is in Half Cheetah, where SMAC achieves an average return of $863.8\pm 38.7$ by the end of training. The next-best algorithm, AC-CG, achieves only $444.6\pm 12.1$, being surpassed by our method by over $1300$ points. Importantly, SMAC also learns three times faster, matching the final performance of AC-CG after only $1.7$ million timesteps. Additionally, note that while SMAC has a higher variance across seeds mid-training, this reduces by the end of learning.
@@
Similar trends can be seen in Hopper, where SMAC reaches the final performance of AC-Adam by using almost three times fewer timesteps. Our method achieves a final average return of $331.2\pm 29.7$, outperforming the second-best performance of AC-Adam with $303.8\pm 27.1$. The stability of SMAC appears to be comparable to that of AC-Adam but lower than that of the other two baselines.
@@
Likewise, in Swimer, SMAC consistently outperforms the second-best baseline, with a final average return of $28.5\pm 0.8$, compared to the second-best return of $23.8\pm 0.9$. Moreover, it learns the fastest, matching the peak performance of AC-CG almost twice as fast. The stability of SMAC is also greater than that of AC-SGD, and competitive with that of AC-Adam and AC-CG.
@@
In Walker, our method learns faster than the baselines in the first $60\%$ of the timesteps. While its final performance of $204.1\pm 87.9$ is lower than those of AC-CG and AC-SGD, which achieve $229.8\pm 23.9$ and $213.4\pm 42.9$, respectively, SMAC reaches the threshold of $115$ (half of the final return of AC-CG) in only $410000$ timesteps, while AC-CG itself needs $39\%$ more. Additionally, our approach outperforms AC-Adam in both average return and stability.
@@
For the Humanoid environment, SMAC achieves the top final performance, $4625\pm 277$ average return, followed closely by AC-SGD with $4539\pm 110$. It greatly outperforms AC-Adam and AC-CG in terms of both final returns and final stability, with only AC-SGD being more stable. While AC-SGD is also slightly faster than our method, SMAC still learns much faster than the other two baselines, matching their final performance after using just $59\%$ of the total allocated training budget.
@@
In Pusher, the performance of SMAC closely follows that of AC-SGD, slightly outperforming the latter with a final average return of $408.2\pm 32.2$. Our method is also more sample efficient, passing the 441-441 return threshold (final performance of AC-CG) $21\%$ faster than AC-CG and $9\%$ faster than AC-SGD. While all agents stabilize by the end of training, only AC-Adam ends up being more stable than SMAC, with AC-SGD having the highest standard deviation.
@@
Finally, we turn our attention to the average log-probabilities of the actions taken during training. The increase in log-probabilities shown by SMAC is stable and correlates with the increase in average returns given in Fig. 3. This may indicate that the policy quickly becomes confident in the actions it chooses, without becoming overconfident. Note that the log-probabilities of our method converge the fastest in $4$ out of $6$ tasks. While the log-probabilities of AC-CG converge first in Humanoid and Pusher, this does not lead to better performance. On the contrary, AC-CG agents become overconfident in their sub-optimal actions in the former task. The slow convergence of the other two agents’ log-probabilities may indicate uncertainty in their actions, even when their performance matches that of SMAC.

### v2

@@ 4.4 SM-ActorCritic
using the Adam optimizer with a learning rate of $\alpha$, where $R_{t}=A_{t}+V_{\phi}(s_{t})$ is the Monte-Carlo return target. The actor is updated by a single natural policy gradient step obtained from the matrix-free Fisher inverse. Since our method is based on Sherman-Morrison formula, we call the resulting algorithm Sherman-Morrison Actor-Critic (SMAC), which is detailed in Algorithm 1.
@@
7:   $g^{k}\leftarrow\nabla_{\theta}\!\log\pi_{\theta}(a\!\mid\!s)A(s,a)$
@@
10:  update critic: $\phi^{k+1}\leftarrow\phi^{k}-\alpha\nabla_{\phi}\sum_{t}(R^{k}_{t}-V_{\phi}(s_{t}))^{2}$
@@ 5 Experimental Setup
To evaluate the performance of our proposed approach, SMAC, we provide experiments in several environments and compare against four baselines. We begin with three simple control environments, before moving on to six MuJoCo tasks (Todorov et al., 2012). For all experiments, we used the Gymnasium framework (Towers et al., 2024).
We implement SMAC on top of the A2C algorithm. In all environments, we evaluate SMAC against three baselines. To measure the benefits of using our NPG approximator, we consider an algorithm that is similar to SMAC in all aspects, except that it uses first-order updates instead of natural gradients. This is equivalent to A2C with a simple stochastic gradient descent (SGD) optimizer, which we denote as AC-SGD. To ensure a fair comparison to vanilla A2C, we also compare against the AC-Adam baseline, where A2C is trained using the more complex Adam optimizer. We also test SMAC against other NPG approximators. The AC-CG baseline uses the conjugate gradient algorithm to directly compute the NPG during the A2C update, without ever computing or storing the Fisher information matrix itself, similarly to TRPO (Schulman et al., 2015a). Finally, the second NPG approximator uses the Kronecker-factored approximate curvature (Martens and Grosse, 2015, KFAC). Our implementation, AC-KFAC, follows the application of KFAC to actor-critic algorithms as proposed by Wu et al. (2017), but omits the trust region.
SMAC and all the baselines we compare against compute advantages using the GAE. This is because GAE handles the bias-variance trade-off better than the simpler advantage estimator used by A2C, and has been shown to be effective in high-dimensional continuous control (Schulman et al., 2015b).
For each of the five algorithms in each environment, we train five models using different random seeds. The results reported throughout training are averaged over these seeds, with standard deviations included. We evaluate agent performance using (undiscounted) returns, averaged across episodes. Besides the maximum average returns achieved, we are also interested in convergence speed.
Furthermore, we also track the average log-probabilities of the actions taken. We treat these as a measure of the agents’ confidence in their actions. We therefore expect a well-performing policy to become more deterministic, achieving both higher log-probabilities and higher returns. On the contrary, we can also diagnose overconfident policies, where an increase in log-probabilities does not correlate to an increase in returns. The agent may therefore prematurely converge to an overly deterministic sub-optimal policy, which is likely to hinder exploration and thus future improvements. We report the log-probabilities measured in the supplementary material.
In all environments, episodes have a horizon of $1000$ timesteps. To increase the readability of the plots, we group timesteps into $100$ bins with no overlap and report the average of each bin. We apply additional smoothing by computing an exponentially weighted mean over the bins, with a smoothing factor of $0.1$. The average performances and standard deviations we report are then computed from these binned and smoothed results. However, for the sake of completeness, we also plot the raw, unbinned, and unsmoothed results in the supplementary material.
AC-SGD AC-Adam AC-CG AC-KFAC SMAC (Ours)
Acrobot $\mathbf{-86.2\pm 4.7}$ $88.4\pm 5.7$ $93.3\pm 11.6$ $170.7\pm 16.5$ $94.1\pm 17.0$
Cartpole $962.3\pm 38.5$ $\mathbf{989.5\pm 14.1}$ $977.9\pm 34.4$ $331.4\pm 142.3$ $973.4\pm 49.5$
Pendulum $2116\pm 678$ $5731\pm 272$ $\mathbf{-190\pm 23}$ $1161\pm 15$ $2226\pm 1138$
Half Cheetah $1672\pm 10.6$ $646\pm 9.7$ $1766\pm 864$ $513\pm 50.3$ $\mathbf{2936\pm 946}$
[…]
Swimmer $12.4\pm 3.0$ $17.3\pm 0.9$ $23.8\pm 1.7$ $24.2\pm 2.5$ $\mathbf{28.5\pm 1.4}$
Walker $213.4\pm 36.9$ $79.3\pm 58.7$ $\mathbf{229.8\pm 38.3}$ $83.7\pm 72.9$ $204.1\pm 58.0$
Humanoid $4539\pm 171$ $3105\pm 345$ $3175\pm 1445$ $260.8\pm 11.15$ $\mathbf{4625\pm 274}$
Pusher $433.6\pm 41.6$ $568.3\pm 16.4$ $441.0\pm 35.6$ $1086.3\pm 16.2$ $\mathbf{-408.2\pm 31.5}$
Table 1: Average returns per episode, measured at the end of training, for each task in both classic control and MuJoCo environments. All results are averaged over 55 seeds ($\pm$ standard deviations).
5.1 Practical Method
@@
The matrix-free updates operate on a single trajectory sample. Although the per-step cost is only $\mathcal{O}(d)$, when applying this per-sample scheme to a full batch of size $b$, the update must be executed b times, so the effective time complexity for the batch becomes $\mathcal{O}(bd)$, so the computational cost can be high to update a whole batch.
To investigate this trade-off, we evaluate an alternative that uses the average gradient of log-probability over all sampled state–action transitions in the batch:
@@
Where $B$ denotes the total number of state-action pairs collected in the batch. The regularized empirical Fisher then becomes
$\tilde{{\bm{F}}}=\lambda\bm{I}+\bar{\ell}^{k}\,(\bar{\ell}^{k})^{\top},$
This is a rank-one moment approximation that assumes vectors in batch have small angular variance. The bias can be written explicitly:
$\mathbb{E}[\tilde{{\bm{F}}}-\hat{{\bm{F}}}]=\frac{1}{B}Cov(\ell)\succeq 0$
so $\tilde{{\bm{F}}}$ underestimates $\hat{{\bm{F}}}$ by a positive semidefinite term that shrinks as gradients concentrate or as $B$ grows.
@@
$\theta^{k+1}=\theta^{k}-\eta\left[\frac{1}{\lambda}g^{k}-\frac{\bar{\ell}^{k}\,(\bar{\ell}^{k})^{\top}g^{k}}{\lambda^{2}+\lambda(\bar{\ell}^{k})^{\top}\bar{\ell}^{k}}\right].$
Results
5.2 Effect of damping coefficient $\lambda$
We use the EF $\hat{{\bm{F}}}^{k}=\lambda{\bm{I}}+\ell^{k}(\ell^{k})^{\top}$ with $\lambda>0$. After applying the Sherman-Morrison formula, its inverse has the closed form
$(\lambda{\bm{I}}+\ell^{k}(\ell^{k})^{\top})^{-1}=\frac{1}{\lambda}({\bm{I}}-\frac{\ell^{k}(\ell^{k})^{\top}}{\lambda+(\ell^{k})^{\top}\ell^{k}}).$
It scales directions unevenly, components perpendicular to $\ell$ are reduced by a factor $\frac{1}{\lambda}$, while the component along $\ell$ is reduced by $\frac{1}{\lambda+(\ell^{k})^{\top}\ell^{k}}$.
Figure 4: Effect of different $\lambda$
@@ Results.
We evaluate SMAC on three Mujoco tasks, compare three damping coeffcient value $\lambda\in\{0.1,1,10\}$. Overall, the results agree with our analysis: too large $\lambda$ slows learning, too small $\lambda$ can make the update noisy, and a moderate $\lambda$ gives the best stability–speed tradeoff on average. The choices in practice are more complex, for example, the choice of $\lambda$ is coupled with the learning rate because both scale the step, so a balance is needed. A feasible path is to use a small damping coefficient to keep the EF term close to its undamped form, and then apply a normalization so the update magnitude is preserved.
5.3 Effect of previous estimate EF
Figure 5: Previous EF used vs. Ours
Using the previous EF not only increases the memory and time complexity to $\mathcal{O}(d^{2})$, but introduces a lag bias: after each update the policy and state distribution change, so the old EF reflects the old distribution and can mis-scale directions when curvature changes quickly. In this experiments we used a mixed estimator ${\bm{F}}^{k}=\beta{\bm{F}}^{k-1}+(1-\ \beta)\ell^{k}(\ell^{k})^{\top}$ with a fixed $\beta=0.5$ to limit reliance on the old EF. Results show that on Swimmer it gives faster early improvement but remains less stable, while on Hopper and Walker it learns more slowly with larger variance. Therefore, if the KL change is strictly constrained, using the previous estimate EF is still feasible, and the choice of $\beta$ should depend on the rate of policy change.
@@ 6 Results
We provide the final performance achieved by training each of the five algorithms on each of the nine tasks. The results in Table 1 show that our method, SMAC, can achieve competitive results in two of the three simple control tasks. More importantly, the benefits of SMAC appear to increase in more complex environments, such as MuJoCo, where SMAC outperforms the other baselines in terms of final returns in all but one task. A more in-depth analysis of these results, with a special focus on convergence rate, is given in this section.
Figure 6: Results on three classic control tasks. We report the average return per episode. All results are first grouped into bins and smoothed, and then
averaged over $5$ random seeds, with shaded areas showing standard deviation.
@@ 6.1 Classic Control
In the classic control environment, our method, SMAC, reaches a high initial average return faster than the baselines on two of the tasks, as shown in Fig. 6 (see Fig. 8 for raw results). This, however, comes at the cost of additional variance across seeds. The overall performance of SMAC is also competitive in those two tasks.
In Acrobot, SMAC only requires $40000$ timesteps to pass the average return threshold of 400-400. This is faster than AC-Adam, AC-SGD, AC-CG, and AC-KFAC, which require $70\%$, $35\%$, $50\%$, and $235\%$ more timesteps, respectively. This accelerated learning is, however, followed by unstable training until approximately $125000$ timesteps. Despite this, SMAC stabilizes and converges, reaching a final performance of $94.1\pm 6.3$, which is similar to three of the baselines, and higher than AC-KFAC.
A similar behavior can be observed in Cartpole. SMAC agents achieve $75\%$ of the maximum return possible in $87000$ timesteps, slightly faster than AC-CG, which needs approximately $10\%$ more timesteps. AC-Adam and AC-SGD are even slower, requiring $28\%$ and $62\%$ more timesteps, respectively, to reach the same threshold. At the same time, the average performance of AC-KFAC never improves above $40\%$ of the optimal policy’s return. SMAC becomes more unstable after this initial period, but stabilizes and reaches a competitive performance of $973.4\pm 29.8$ by the end of training.
In contrast, in the last environment, Pendulum, our method fails to outperform AC-CG in terms of convergence speed and overall performance. AC-CG requires only $1.5$ million timesteps to match the final performance of SMAC, indicating that our proposed approximation of the FIM can struggle to improve learning in some environments. Moreover, SMAC is also outperformed by AC-KFAC, which quickly finds a good policy but then stops learning. However, SMAC still achieves a final return of $2226\pm 859$, outperforming AC-Adam and matching AC-SGD, with the former also displaying much slower and unstable learning.
As proposed in Sec. 5, the validity of our results is also confirmed by the action log-probabilities recorded (Fig. 10). In Acrobot and Cartpole, the two tasks in which SMAC performs competitively, the log-probabilities converge faster than those of the baseline methods. Together with the increase in performance achieved by SMAC in these tasks, this may indicate that our agents can rapidly become confident in their policies when taking actions that yield increasingly higher returns. At the same time, in Pendulum, the log-probabilities of SMAC match those of AC-SGD and are bounded above by AC-CG. This could indicate that our method becomes overconfident in its actions too early, finally leading to a sub-optimal policy.
@@ 6.2 MuJoCo
In the more complex MuJoCo environment, we have found that throughout training, our method, SMAC, consistently outperforms or matches the baseline algorithms in terms of performance in $5$ out of $6$ tasks, as shown in Fig. 7 (see Fig. 9 for raw results). In the sixth task, SMAC still achieves comparable results.
Figure 7: Results on six MuJoCo tasks. We report the average return per episode. All results are first grouped into bins and smoothed, and then
@@
averaged over $5$ random seeds, with shaded areas showing standard deviation.
SMAC shows improvements over existing methods in Half Cheetah, where it achieves an average return of $2936\pm 946$, surpassing the next-best algorithm, AC-CG, by over $1100$ points. While it learns more slowly for the first $2$ million timesteps, SMAC ultimately matches the final performance of AC-CG after only $3.4$ million timesteps, which represents just $68\%$ of the training budget. It then continues to improve, while AC-CG plateaus. This superior performance comes at the cost of lower stability, with both SMAC and AC-CG being much more unstable than AC-KFAC, and the simpler AC-Adam and AC-SGD.
@@
Similar trends can be observed in Hopper, where SMAC achieves the final performance of AC-Adam using almost three times fewer timesteps. Our method achieves a final average return of $331.2\pm 29.7$, outperforming the second-best performances of AC-Adam and AC-KFAC by approximately $30$ points. The stability of SMAC appears to be comparable to that of AC-Adam, higher than that of AC-KFAC, but lower than that of the other two baselines.
@@
Likewise, in Swimer, SMAC consistently outperforms the second-best baselines, with a final average return of $28.5\pm 0.8$, which is approximately $5$ points higher than those of AC-CG and AC-KFAC. Moreover, it learns the fastest, matching the peak performance of AC-CG almost twice as fast. The stability of SMAC is also greater than that of AC-SGD, AC-CG, and AC-KFAC, and competitive with that of AC-Adam.
@@
In Walker, our method learns faster than the baselines in the first $60\%$ of the timesteps. While its final performance of $204.1\pm 87.9$ is lower than those of AC-CG and AC-SGD, which achieve $229.8\pm 23.9$ and $213.4\pm 42.9$, respectively, SMAC reaches the threshold of $115$ (half of the final return of AC-CG) in only $410000$ timesteps, while AC-CG itself needs $39\%$ more. Additionally, our approach outperforms AC-Adam and AC-KFAC in both average return and stability.
@@
For the Humanoid environment, SMAC achieves the top final performance of $4625\pm 277$ average return, followed closely by AC-SGD with $4539\pm 110$. It greatly outperforms AC-Adam and AC-CG in terms of both final returns and final stability, with only AC-SGD and AC-KFAC being more stable. However, the performance of AC-KFAC is much lower than that of any of the other algorithms. While AC-SGD is slightly faster than our method, SMAC still learns much faster than AC-Adam and AC-CG, matching their final performance after using just $59\%$ of the total allocated training budget.
@@
In Pusher, the performance of SMAC closely follows that of AC-SGD, slightly outperforming the latter with a final average return of $408.2\pm 32.2$. Our method is also more sample efficient, passing the 441-441 return threshold (final performance of AC-CG) $21\%$ faster than AC-CG and $9\%$ faster than AC-SGD. While all agents stabilize by the end of training, only AC-Adam and AC-KFAC end up being more stable than SMAC, with AC-SGD having the highest standard deviation.
@@
Finally, we turn our attention to the average log-probabilities of the actions taken during training (Fig. 11), as proposed in Sec. 5. The increase in log-probabilities shown by SMAC is stable and correlates with the increase in average returns given in Fig. 7. This may indicate that the policy quickly becomes confident in the actions it chooses, without becoming overconfident. Note that the log-probabilities of our method converge the fastest in $2$ out of $6$ tasks. While the log-probabilities of AC-CG converge first in Humanoid and Pusher, this does not lead to better performance. On the contrary, in the former task, AC-CG agents become overconfident in their sub-optimal actions. While in Half Cheetah the log-probabilities of AC-CG converge first, SMAC ultimately achieves a higher confidence in its actions, which matches its superior final performance. A similar trend appears in Walker, with AC-KFAC converging the fastest, but to a sub-optimal policy. The slow convergence of the other two agents’ log-probabilities in most tasks may indicate uncertainty in their actions, even when their performance matches that of SMAC.